OpenStudy (hocuspucus292333393):

Use absolute value to find the distance between point A and point B.

OpenStudy (hocuspucus292333393):

@Joshiscool help

jimthompson5910 (jim_thompson5910):

B is the first letter shown in B-A so the value where B is located will go in that first box

OpenStudy (hocuspucus292333393):

THIS

jimthompson5910 (jim_thompson5910):

you use the number, not the letter

jimthompson5910 (jim_thompson5910):

where is B on the number line?

OpenStudy (hocuspucus292333393):

7

jimthompson5910 (jim_thompson5910):

yes so 7 goes in that first box

jimthompson5910 (jim_thompson5910):

where is A located?

OpenStudy (hocuspucus292333393):

-6

jimthompson5910 (jim_thompson5910):

-6 goes in that second box

OpenStudy (hocuspucus292333393):

LIKE THIS

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (hocuspucus292333393):

SO the last one bottom

jimthompson5910 (jim_thompson5910):

find the distance from B to A, or from A to B

jimthompson5910 (jim_thompson5910):

you can count the spaces or evaluate 7-(-6)

jimthompson5910 (jim_thompson5910):

hint: the negatives cancel to form a positive similar example: 3-(-5) = 3+5 = 8

OpenStudy (hocuspucus292333393):

so am I right

OpenStudy (vuriffy):

@HocusPucus292333393 That was an example, not the answer..

OpenStudy (vuriffy):

What's 7 + 6?

OpenStudy (hocuspucus292333393):

its 13

OpenStudy (vuriffy):

Yes, that's your answer. 7-(-6) = 13.

OpenStudy (hocuspucus292333393):

this

jimthompson5910 (jim_thompson5910):

yes if you count the spaces, you should count 13 spaces

OpenStudy (vuriffy):

Yes!^

OpenStudy (hocuspucus292333393):

can u guys help with 5 easy questions